A researcher claims that 10-year old children watch 6.6 hours of TV daily on average. You try to verify this for a sample of size 100 with mean 6.1 hours and standard deviation 2.5 hours. 


1
Expert's answer
2021-07-13T05:17:42-0400

The following null and alternative hypotheses need to be tested:

"H_0:\mu=6.6"

"H_1:\mu\not=6.6"

This corresponds to a two-tailed test, for which a t-test for one mean, with unknown population standard deviation, using the sample standard deviation, will be used.

Based on the information provided, the significance level is "\alpha=0.05," "df=n-1=100-1=99" degrees of freedom and the critical value for a two-tailed test is "t_c=1.9842."

The rejection region for this two-tailed test is "R=\{t: |t|>1.9842\}"

The t-statistic is computed as follows:


"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{6.1-6.6}{2.5/\sqrt{100}}=-2"

Since it is observed that "|t|=2>1.9842=t_c," it is then concluded that the null hypothesis is rejected.


Using the P-value approach: The p-value for "\alpha=0.05, df=99" degrees of freedom, two-tailed, "t=-2" is "p=0.04824," and since "p=0.04824<0.05=\alpha," it is concluded that the null hypothesis is rejected.

Therefore, there is enough evidence to claim that the population mean "\mu"  is different than "6.6," at the "\alpha=0.05" significance level.